# How weak are weak factors? Uniform inference for signal strength in signal plus noise models

### v1

@@ 5.2. S&P100
Analyzing stock returns is essential for understanding market dynamics, evaluating investment performance, and guiding both individual and institutional investment strategies. A key statistical object in this context is the covariance matrix of stock returns, which plays a central role in portfolio optimization, such as in the Markowitz mean–variance framework. The vast “factor zoo” – the large number of potential variables proposed to explain stock returns – highlights the practical challenge of distinguishing meaningful factors from noise in high-dimensional settings (see Cochrane (2011) for an influential discussion). Here we demonstrate how our methodology can be applied to the sample covariance matrix of weekly S&\&&P100100100100 stock returns. We use data from Bykhovskaya and Gorin (2022), which covers 92 stocks over the period from January 1, 2010, to January 1, 2020.
Figure 4 presents the full spectrum of the sample covariance matrix and identifies ten eigenvalues – 0.0365, 0.005, 0.004, 0.003, 0.0026, 0.0021, 0.0018, 0.0015, 0.00139, 0.001360.03650.0050.0040.0030.00260.00210.00180.00150.001390.001360.0365,\,0.005,\,0.004,\,0.003,\,0.0026,\,0.0021,\,0.0018,\,0.0015,\,0.00139,\,0.001360.0365 , 0.005 , 0.004 , 0.003 , 0.0026 , 0.0021 , 0.0018 , 0.0015 , 0.00139 , 0.00136 – that lie to the right of the theoretical Marchenko–Pastur upper edge, $\lambda_{+}=0.0013$. To better fit the empirical data, we adopt an effective parameter value $\gamma^{2}=0.4$, in contrast to the true value $N/S=0.18$. This adjustment may reflect temporal dependence in the data, which effectively reduces the sample size and increases $\gamma^{2}=0.4$. We also set $\sigma=0.02$ to align the overall variance of the eigenvalues in the data.
@@
Figure 4. S&\&&P100100100100 sample covariance: signals, their $95\%$ confidence intervals, noise, and Marchenko-Pastur distribution with $N=92$, $\gamma^{2}=0.4$, $\sigma=0.02$.
@@
Financial stocks are typically classified into cyclical and non-cyclical (defensive) categories, depending on whether their performance tends to track economic business cycles. These groups are generally assumed to be uncorrelated, aside from exposure to a common “market” factor. Bykhovskaya and Gorin (2025) identify three non-zero canonical correlations between these two groups, suggesting the presence of three common factors. Here we revisit their analysis using the same data set to assess whether these observed correlations reflect genuine signals or could instead be attributed to noise.
@@ 6.1. Non-Gaussian noise
The four models in Section 2 are all based on Gaussian noise matrices. A natural generalization is to replace the Gaussian vectors with more general random vectors having the same mean and covariance. It is well known (see Lee and Yin (2014); Ding and Yang (2018); Yang (2022a) and the broader reviews Deift and Gioev (2009); Tao and Vu (2012); Erdős and Yau (2017)) that for pure noise models without any signal, the distribution of the largest eigenvalues remains unchanged in many non-Gaussian settings. This raises the natural question of whether the same robustness extends to the confidence intervals discussed in Section 3.
Figure 6 shows the results of Monte Carlo simulations of confidence intervals for Wigner matrices with a single spike and non-Gaussian noise. These can be compared to the Gaussian case in Figure 1(a). The bold gray line depicts sample confidence intervals based on $10^{5}$ simulations of $100\times 100$ matrices. The noise matrix $\mathcal{E}$ in (2.1) is still defined as $\mathcal{E}=\frac{1}{\sqrt{2N}}(\mathcal{Z}+\mathcal{Z}^{\mathsf{T}})$, with $\mathcal{Z}$ having i.i.d. entries. We vary the distribution of these entries, each rescaled to have mean 00 and variance $1$, to be: (i) uniform on $[0,1]$, (ii) Bernoulli with success probability $p=1/2$, or (iii) Binomial with parameters $n=2$, $p=0.3$. We also consider two signal vectors: a localized signal $\mathbf{u}^{*}=(1,0,\dots,0)^{\mathsf{T}}$ and a delocalized signal $\mathbf{u}^{*}=\tfrac{1}{\sqrt{N}}(1,1,\dots,1)^{\mathsf{T}}$. While in the Gaussian setting the choice of $\mathbf{u}^{*}$ is irrelevant due to rotational invariance, this is no longer true for general noise distributions.
Figure 6 shows that while the sample confidence intervals remain reasonably close to those constructed via the procedure in Section 3, the agreement is notably better for delocalized signals. For the localized signal significant deviations appear at larger values of the observed eigenvalue $\lambda_{1}$. A theoretical explanation follows directly from the model $\mathbf{A}=\theta\cdot\mathbf{u}^{*}(\mathbf{u}^{*})^{\mathsf{T}}+\mathcal{E}$ in (2.1). When $\mathbf{u}^{*}=(1,0,\dots,0)^{\mathsf{T}}$, the parameter $\theta$ enters $\mathbf{A}$ only through its sum with the $(1,1)$ entry of $\mathcal{E}$, so the distribution of that single element directly affects any estimate of $\theta$. For $\theta>\theta^{c}$, this effect was rigorously analyzed in Capitaine et al. (2009). By contrast, when $\mathbf{u}^{*}=\tfrac{1}{\sqrt{N}}(1,1,\dots,1)^{\mathsf{T}}$, the projection of the noise onto the signal direction aggregates many independent entries. Thus, by the central limit theorem, the influence of the individual noise distribution diminishes as $N$ grows, leading to behavior indistinguishable from the Gaussian case. We expect similar robustness for other delocalized signals, though a precise definition of “delocalized” is still lacking in the literature. One rigorous result in this direction is given in Benaych-Georges et al. (2011), which shows that if the entries of $\mathbf{u}^{*}$ are i.i.d. and satisfy certain technical conditions, then for $\theta>\theta^{c}$, the asymptotic distribution of $\lambda_{1}$ is indeed independent of whether the noise is Gaussian.
@@ 6.2. General models and empirical distributions of eigenvalues
In all four models from Section 2 the empirical distributions of eigenvalues take specific parametric forms, as summarized in Table 3. Although our estimation procedure relies only on the largest eigenvalues, the formulas in Table 4 were derived using information from the empirical distributions in Table 3.
In other signal plus noise models and some empirical data sets the eigenvalue histograms may differ substantially from the four cases we consider. For $\theta>\theta^{c}$, the fluctuations of the largest eigenvalues have been analyzed in various alternative settings; see, e.g., Benaych-Georges et al. (2011); Benaych-Georges and Nadakuditi (2012); Onatski (2012). These works develop analogues of (2.3), (2.9), and (2.13), but the formulas for $\theta^{c}$, $\lambda_{+}$, $\lambda(\theta)$, and $V(\theta)$ become more complex. We conjecture that the confidence interval procedure from Section 3 remains valid in such broader contexts, with only the model-dependent parameters from Table 4 needing to be updated.
One applied setting of particular interest is the approximate factor model, which resembles the setup in Section 2.3 but allows the noise matrix $\mathcal{E}$ to have a more complex correlation structure rather than being i.i.d. To apply our confidence interval procedure in this context, one can proceed as follows: first, select a value for $\lambda_{+}$; then use all eigenvalues below $\lambda_{+}$ to estimate the empirical distribution of noise eigenvalues (replacing the parametric forms in Table 3); next, substitute this estimate into the formulas from Onatski (2012) for $\lambda(\theta)$ and $V(\theta)$; and finally, apply our method to construct confidence intervals for the eigenvalues exceeding $\lambda_{+}$. Choosing $\lambda_{+}$ optimally is delicate; one approach, suggested in Onatski (2010, Section 4), is to select it based on the characteristic $\sqrt{x}$ behavior of the eigenvalue density near the edge – a feature clearly visible in the four models of Table 3 and present in many other cases.
@@ 7. Conclusion
Our procedure is robust to both weak and critical signals, enabling practitioners to distinguish between informative and non-informative components without imposing assumptions on signal strength. Moreover, our methodology reveals a surprising universality: despite differences in the statistical structure of the models, the same transition process governs the fluctuations of their top eigenvalues. We believe that this suggests deeper underlying principles in high-dimensional inference and opens a broad avenue for future research in more general signal plus noise settings.
8. Appendix A: Random Stieltjes transform at the edge and general asymptotic theorem
Our goal in this section is to rigorously introduce the Airy–Green function $\mathcal{G}(w)$ and limit theorems related to it. Universally, in all settings we study, the equations connecting the parameters of the spiked model with observed largest eigenvalues will be asymptotically written in terms of $\mathcal{G}(w)$.
[…]
$\sup_{z\in W}d(f_{n}(z),f(z))\to 0\leavevmode\nobreak\ ,$
where
$d(\zeta,\zeta^{\prime})=\frac{|\zeta-\zeta^{\prime}|}{\sqrt{1+|\zeta|^{2}}\sqrt{1+|\zeta^{\prime}|^{2}}}$
[…]
@@ 8.1. The function $\mathcal{G}(w)$
In this subsection we restate, elaborate on, and prove the statements from Section 4.1.
@@
The conditional expectation $\mathbb{E}(\mathcal{G}_{\beta}(w)|\{\mathfrak{a}_{j}\})$ (equivalent to replacing $\frac{1}{\beta}\xi^{2}_{j}$ by $1$) was also used recently in another context by Huang and Zhang (2024).
[…]
@@
(8.2) $\lim_{\begin{smallmatrix}w\to\infty\\ \mathrm{Re}(w)\geq 0\end{smallmatrix}}\left|\mathcal{G}(w)+\sqrt{w}\right|=0,\qquad\text{almost surely.}$.
@@
Let $\rho(x)\mathrm{d}x$ be the first correlation measure of $\{\mathfrak{a}_{j}\}$, which means that for any compactly supported bounded $f(x)$ with finitely many discontinuity points, we have
$\mathbb{E}\sum_{j=1}^{\infty}f(\mathfrak{a}_{j})=\int_{-\infty}^{\infty}f(x)\rho(x)\mathrm{d}x.$
Then:
@@
    At $\infty$- ∞, $\rho(x)$ has the asymptotics
    (8.5) $\rho(x)=\frac{(-x)^{1/2}}{\pi}+O(1/|x|),\quad x\to-\infty\leavevmode\nobreak\ .$
@@
(8.13) $\mathbb{E}\bigl{(}\#\{j\geq 1\mid\mathfrak{a}_{j}>-T\}\bigr{)}=\frac{2}{3\pi}T^{3/2}+O(\ln(T)),\quad T\to+\infty\leavevmode\nobreak\ .$
@@
For $\mathfrak{a}_{1}>\mathfrak{a}_{2}>\mathfrak{a}_{3}>\dots$ being the Airy1 point process,
(8.15) $\lim_{T\to+\infty}\frac{\mathrm{Var}\bigl{(}\#\{j\geq 1\mid\mathfrak{a}_{j}>-T\}\bigr{)}}{\ln(T)}=\frac{11}{6\pi^{2}}.$
@@
where $\{\mathfrak{a}_{i}\}$ and $\{\mathfrak{a}^{\prime}_{i}\}$ are two independent copies of the Airy1 point process, $\mathfrak{a}^{\beta=2}_{i}$ is the Airy2 point process, and “even” is the operation of removing all particles with odd indices, i.e., keeping the 2nd, 4th, 6th, etc, largest ones. From Soshnikov (2000, Theorem 1), it is known that as $T\to+\infty$, $\mathrm{Var}\bigl{(}\#\{j\geq 1\mid\mathfrak{a}_{j}^{\beta=2}>-T\}\bigr{)}\sim\frac{11}{12\pi^{2}}\ln(T)$. Through (8.16) this implies
$\lim_{T\to+\infty}\frac{\mathrm{Var}\bigl{(}\#\{j\geq 1\mid\mathfrak{a}_{j}>-T\}+\#\{j\geq 1\mid\mathfrak{a}^{\prime}_{j}>-T\}\bigr{)}}{\ln(T)}=4\cdot\frac{11}{12\pi^{2}}=\frac{11}{3\pi^{2}},$
because the “even” operator divides the number of particles by two (up to error of at most $1$), and therefore divides the variance by four (up to an error negligible as $T\to\infty$). Since variances for independent random variables are added, we get (8.15). ∎
@@
For each $\varepsilon>0$ there exists a random variable $\mathfrak{J}=\mathfrak{J}(\varepsilon)$, such that almost surely we have
(8.17) $\left|\mathfrak{a}_{j}+\left(\frac{3\pi j}{2}\right)^{2/3}\right|\leq j^{\varepsilon},\qquad\text{for all }j>\mathfrak{J}.$
@@
|#$\bigr{|}\#\{j\geq 1\mid\mathfrak{a}_{j}>-n\}-\mathbb{E}\#\{j\geq 1\mid\mathfrak{a}_{j}>-n\}\bigl{|}>n^{1/2+\varepsilon}.$
@@
|#$\bigr{|}\#\{j\geq 1\mid\mathfrak{a}_{j}>-n\}-\mathbb{E}\#\{j\geq 1\mid\mathfrak{a}_{j}>-n\}\bigl{|}\leq n^{1/2+\varepsilon},\qquad\text{for all }n>\mathfrak{n}.$
@@
|#$\left|\#\{j\geq 1\mid\mathfrak{a}_{j}>-n\}-\frac{2}{3\pi}n^{3/2}\right|\leq n^{1/2+\varepsilon},\qquad\text{for all }n>\mathfrak{n}.$
@@
Note that
$\frac{1}{w-\mathfrak{a}_{j}}-\frac{1}{\mathbf{i}-\mathfrak{a}_{j}}=\frac{\mathbf{i}-w}{(w-\mathfrak{a}_{j})(\mathbf{i}-\mathfrak{a}_{j})}.$
Hence, using (8.17), the $j$–th term in the first sum of (8.20) decays as $j^{-4/3}$ and the sum is absolutely convergent, uniformly in $w$ bounded away from $\mathfrak{a}_{j}$. For the second sum, its imaginary part is
@@
Using (8.17), for large $j$ we have $-\mathfrak{a}_{j}=\left(\frac{3\pi j}{2}\right)^{2/3}+O(j^{\varepsilon})$, and, therefore, for large $m$ and $n$:
@@
We check the Cauchy criterion for (8.21) and compute the difference of its values at $x=-y$ and $x=-z$ for large $y>z>0$. Using (8.22), we get
@@
where $n(y)$ is the index $j$ for the closest to $-y$ point $\mathfrak{a}_{j}$ and $m(z)$ is the index for the closest to $-z$ point $\mathfrak{a}_{j}$. Assuming $y$ and $z$ large enough, so that $n(y)>\mathfrak{J}$ and $m(z)>\mathfrak{J}$, we can use (8.17) and get
@@
We condition on $\{\mathfrak{a}_{j}\}$ and note that we deal with a sum of independent mean 00 random variables. Hence, by the Kolmogorov two-series theorem (see, e.g., Durrett (2019, Theorem 2.5.6)), the almost sure convergence would follow from the convergence of the sum of (conditional) variances, i.e., convergence of the series
@@
$\mathcal{G}(w)\\ =\lim_{x\to-\infty}\left[\sum_{j:\,\mathfrak{a}_{j}>x}\left(\frac{1}{w+\left(\frac{3\pi j}{2}\right)^{2/3}}+\frac{\xi_{j}^{2}-1}{w+\left(\frac{3\pi j}{2}\right)^{2/3}}+\xi_{j}^{2}\left[\frac{\left(\frac{3\pi j}{2}\right)^{2/3}+\mathfrak{a}_{j}}{(w-\mathfrak{a}_{j})(w+\left(\frac{3\pi j}{2}\right)^{2/3})}\right]\right)-\frac{2}{\pi}\sqrt{-x}\right].$
@@
Let us show that as $w$ becomes large (constrained by $\mathrm{Re}(w)\geq 0$), the second and third sums in (8.25) tend to 00.

### v2

@@ 5.2. S&P100
Analyzing stock returns is essential for understanding market dynamics, evaluating investment performance, and guiding both individual and institutional investment strategies. A key statistical object in this context is the covariance matrix of stock returns, which plays a central role in portfolio optimization, such as in the Markowitz mean–variance framework. The vast “factor zoo” – the large number of potential variables proposed to explain stock returns – highlights the practical challenge of distinguishing meaningful factors from noise in high-dimensional settings (see Cochrane (2011) for an influential discussion). Here we demonstrate how our methodology can be applied to the sample covariance matrix of weekly S&\&P100100 stock returns. We use data from Bykhovskaya and Gorin (2022), which covers 92 stocks over the period from January 1, 2010, to January 1, 2020.
Before turning to the empirical spectrum, it is important to address a potential concern regarding the applicability of PCA-based methods in financial data. Financial returns may exhibit heavy tails, complicating factor identification via PCA. For pure-noise covariance matrices, a sharp transition occurs at tail index $\alpha=4$ (Yin et al., 1988; Bai et al., 1988): lighter tails keep the largest eigenvalue at the Marchenko–Pastur edge, while heavier tails can produce spurious spikes. Although short-horizon returns often have $\alpha\approx 3$, weekly and longer-horizon returns typically have $\alpha\geq 5$ (Gopikrishnan et al., 1999; Gabaix, 2009; Fan and Yao, 2017), as aggregation and the central limit theorem dampen extremes. This places weekly returns in a regime where spurious tail-driven spikes are less likely to arise.
@@
Figure 4. S&\&P100100 sample covariance: signals, their $95\%$ confidence intervals, noise, and Marchenko-Pastur distribution with $N=92$, $\gamma^{2}=0.4$, $\sigma=0.02$.
Confidence Intervals $\times 10^{4}$
$10^{4}\lambda_{i}$ $10^{4}\theta_{i}$ Our method Gaussian method $\sharp 1$ Gaussian method $\sharp 2$
365 358 [347  369] [347  369] [347  369]
52.4 45 [41.5  49.7] [41.3  49.5] [41.4  49.7]
41.7 35 [31.3  38.5] [31.0  38.3] [31.2  38.5]
30.7 23 [20.7  26.9] [20.4  26.5] [20.6  26.7]
26.1 19 [16.2  21.9] [16.0  21.6] [16.2  21.8]
20.7 13 [10.9  15.8] [10.7  15.6] [10.9  15.8]
17.6 10 [7.84  12.4] [7.63  12.1] [7.83  12.3]
15.1 7.0 [5.04  8.90] [4.86  9.07] [5.10  9.27]
13.9 5.3 [3.11  7.65] [3.12  7.39] [3.63  7.59]
13.6 4.8 [2.24  7.26] [2.57  6.98] [3.35  7.17]
Table 6. S&P100: Confidence intervals $\times 10^{4}$ based on different methods.
[…]
@@
Financial stocks are typically classified into cyclical and non-cyclical (defensive) categories, depending on whether their performance tracks economic business cycles. These groups are generally assumed to be uncorrelated, aside from exposure to a common “market” factor. Bykhovskaya and Gorin (2025) identify three non-zero canonical correlations between these two groups, suggesting the presence of three common factors. Here we revisit their analysis using the same data set to assess whether these observed correlations reflect genuine signals or could instead be attributed to noise.
@@ 6.1. Non-Gaussian noise
The four models in Section 2 are based on Gaussian noise matrices. A natural generalization is to replace the Gaussian vectors with more general random vectors having the same mean and covariance. It is well known (see Lee and Yin (2014); Ding and Yang (2018); Yang (2022a) and the broader reviews Deift and Gioev (2009); Tao and Vu (2012); Erdős and Yau (2017)) that for pure noise models without signals, the distribution of the largest eigenvalues remains unchanged in many non-Gaussian settings. This raises the question of whether the robustness extends to the confidence intervals of Section 3.
@@
Figure 6 shows the results of Monte Carlo simulations of confidence intervals for Wigner matrices with a single spike and non-Gaussian noise. These can be compared to the Gaussian case in Figure 1(a). The thick gray line depicts sample confidence intervals based on $10^{5}$ simulations of $100\times 100$ matrices. The noise matrix $\mathcal{E}$ in (2.1) is still defined as $\mathcal{E}=\frac{1}{\sqrt{2N}}(\mathcal{Z}+\mathcal{Z}^{\mathsf{T}})$, with $\mathcal{Z}$ having i.i.d. entries. We vary the distribution of these entries, each rescaled to have mean 0 and variance $1$, to be: (i) uniform on $[0,1]$, (ii) Bernoulli with success probability $p=1/2$, or (iii) Binomial with parameters $n=2$, $p=0.3$. We also consider two signal vectors: a localized signal $\mathbf{u}^{*}=(1,0,\dots,0)^{\mathsf{T}}$ and a delocalized signal $\mathbf{u}^{*}=\tfrac{1}{\sqrt{N}}(1,1,\dots,1)^{\mathsf{T}}$. While in the Gaussian setting the choice of $\mathbf{u}^{*}$ is irrelevant due to rotational invariance, this is no longer true for general noise distributions.
Figure 6 shows that, while the sample confidence intervals remain reasonably close to those constructed via the procedure in Section 3, the agreement is notably better for delocalized signals. For the localized signal significant deviations appear at larger values of the observed eigenvalue $\lambda_{1}$. A heuristic explanation follows directly from the model $\mathbf{A}=\theta\cdot\mathbf{u}^{*}(\mathbf{u}^{*})^{\mathsf{T}}+\mathcal{E}$ in (2.1). When $\mathbf{u}^{*}=(1,0,\dots,0)^{\mathsf{T}}$, the parameter $\theta$ enters $\mathbf{A}$ only through its sum with the $(1,1)$ entry of $\mathcal{E}$, so the distribution of that single element directly affects any estimate of $\theta$. By contrast, when $\mathbf{u}^{*}=\tfrac{1}{\sqrt{N}}(1,1,\dots,1)^{\mathsf{T}}$, the projection of the noise onto the signal direction aggregates many independent entries. Thus, by the central limit theorem, the influence of the individual noise distribution diminishes as $N$ grows, leading to behavior indistinguishable from the Gaussian case. Similar robustness is expected for other delocalized signals.
For spiked Wigner matrices with supercritical signals ($\theta>\theta^{c}$), non-Gaussian cases have been rigorously analyzed in several papers. For localized signals, non-trivial dependence of the asymptotics of $\lambda_{1}$ on the distribution of the noise has been established in Capitaine et al. (2009, 2012); Pizzo et al. (2013); Knowles and Yin (2013, 2014). In contrast, universality of the limit for delocalized signals has been shown under various conditions (including different definitions of “delocalized”) in Féral and Péché (2007); Benaych-Georges et al. (2011); Capitaine et al. (2012); Pizzo et al. (2013); Renfrew and Soshnikov (2013); Knowles and Yin (2013, 2014). Also Knowles and Yin (2013, Remark 2.18) explain that even for localized signals the dependence of the limit on the noise distribution is washed out as $\theta$ approaches $\theta^{c}$. Similar phenomenology is expected (and in some cases proven) for other signal plus noise models.
@@ 6.2. General models and empirical distributions of eigenvalues
In all four models from Section 2 the empirical eigenvalue distributions take specific parametric forms, as summarized in Table 3. Although our estimation procedure relies only on the largest eigenvalues, the formulas in Table 4 were derived using empirical distributions in Table 3.
In other signal plus noise models and some empirical data sets the eigenvalue histograms may differ substantially from the four cases we consider. For $\theta>\theta^{c}$, the fluctuations of the largest eigenvalues have been analyzed in alternative settings; see, e.g., Benaych-Georges et al. (2011); Benaych-Georges and Nadakuditi (2012); Onatski (2012). These works develop analogues of (2.3), (2.9), and (2.13), but the formulas for $\theta^{c}$, $\lambda_{+}$, $\lambda(\theta)$, and $V(\theta)$ become more complex. We conjecture that the confidence interval procedure from Section 3.2 remains valid in such broader contexts, with updated model-dependent parameters from Table 4.
One applied setting of particular interest is the approximate factor model, which resembles the setup in Section 2.3 but allows the noise matrix $\mathcal{E}$ to have a more complex correlation structure rather than being i.i.d. To apply our confidence interval procedure here, proceed as follows: first, select a value for $\lambda_{+}$; then use all eigenvalues below $\lambda_{+}$ to estimate the empirical distribution of noise eigenvalues (replacing the parametric forms in Table 3); next, substitute this estimate into the formulas from Onatski (2012) for $\lambda(\theta)$ and $V(\theta)$; and finally, apply our method to construct confidence intervals for the eigenvalues exceeding $\lambda_{+}$. Choosing $\lambda_{+}$ optimally is delicate; one approach, suggested in Onatski (2010, Section 4), is to select it based on the characteristic $\sqrt{x}$ behavior of the eigenvalue density near the edge – a feature clearly visible in the four models of Table 3 and present in many other cases.
@@ 7. Conclusion
Our procedure is robust to both weak and critical signals, enabling practitioners to distinguish between informative and non-informative components without imposing assumptions on signal strength. Our methodology reveals a surprising universality: despite differences in the statistical structure of the models, the same transition process governs the fluctuations of their top eigenvalues. This suggests deeper underlying principles in high-dimensional inference and opens an avenue for future research in more general signal plus noise settings.
8. Appendix A: Random Stieltjes transform at the edge and asymptotics
Our goal in this section is to rigorously introduce the Airy–Green function $\mathcal{G}(w)$ and related limit theorems. In all settings we study, the equations linking the spiked model parameters to the observed largest eigenvalues can be asymptotically expressed via $\mathcal{G}(w)$.
We first recall the definitions of meromorphic functions and their convergence for use in the proofs. A function $f(z)$ is meromorphic if it is analytic on $\mathbb{C}\setminus\{x_{j}\}$, where $\{x_{j}\}$ is at most a countable set of isolated poles of finite order. Equivalently, $f$ is analytic from $\mathbb{C}$ to the Riemann sphere $\overline{\mathbb{C}}$. For meromorphic $f_{n}$ and $f$, we write $f_{n}\overset{\text{mer}}{\rightarrow}f$ if for any compact $W\subset\mathbb{C}$ we have
$\sup_{z\in W}d(f_{n}(z),f(z))\to 0\penalty 10000\ ,\qquad\text{ with }\qquad d(\zeta,\zeta^{\prime})=\frac{|\zeta-\zeta^{\prime}|}{\sqrt{1+|\zeta|^{2}}\sqrt{1+|\zeta^{\prime}|^{2}}},$
the spherical distance. Equivalently, $f_{n}\overset{\text{mer}}{\rightarrow}f$ if and only if $f_{n}\to f$ uniformly on any compact $W\subset\mathbb{C}$ not containing the poles of $f$.
@@ 8.1. The function $\mathcal{G}(w)$
We restate and prove the statements from Section 4.1.
@@
The conditional expectation $\mathbb{E}(\mathcal{G}(w)|\{\mathfrak{a}_{j}\})$ (equivalent to replacing $\xi^{2}_{j}$ by 11) was used recently in another context by Huang and Zhang (2024). The restriction of $\mathcal{G}(w)$ to real $w>\mathfrak{a}_{1}$ resembles the objects studied in Baik et al. (2021).
The theorem asserts that almost surely the functions converge in the topology $\overset{\text{mer}}{\rightarrow}$.
@@
(8.2) $\lim_{\begin{smallmatrix}w\to\infty\\ \mathrm{Re}(w)\geq 0\end{smallmatrix}}\left|\mathcal{G}(w)+\sqrt{w}\right|=0,\qquad\text{almost surely.}$
@@
Let $\rho(x)\mathrm{d}x$ be the first correlation measure of $\{\mathfrak{a}_{j}\}$, which means that for any compactly supported bounded $f(x)$ with finitely many discontinuity points, we have $\mathbb{E}\sum_{j=1}^{\infty}f(\mathfrak{a}_{j})=\int_{-\infty}^{\infty}f(x)\rho(x)\mathrm{d}x.$ Then:
@@
    As $x\to$-$\infty$, $\rho(x)$ has the asymptotics $\rho(x)=\frac{(-x)^{1/2}}{\pi}+O(1/|x|).$
@@
(8.12) $\mathbb{E}\bigl(\#\{j\geq 1\mid\mathfrak{a}_{j}>-T\}\bigr)=\frac{2}{3\pi}T^{3/2}+O(\ln(T)),\quad T\to+\infty\penalty 10000\ .$
@@
For $\mathfrak{a}_{1}>\mathfrak{a}_{2}>\mathfrak{a}_{3}>\dots$ being the Airy1 point process, there exists a constant $C_{1}>0$ such that
(8.14) $\lim_{T\to+\infty}\frac{\mathrm{Var}\bigl(\#\{j\geq 1\mid\mathfrak{a}_{j}>-T\}\bigr)}{\ln(T)}=\frac{3}{2\pi^{2}}.$
@@
where $\{\mathfrak{a}_{i}\}$ and $\{\mathfrak{a}^{\prime}_{i}\}$ are two independent copies of the Airy1 point process, $\mathfrak{a}^{\beta=2}_{i}$ is the Airy2 point process, and “even” is the operation of removing all particles with odd indices, i.e., keeping the 2nd, 4th, 6th, etc, largest ones. From Soshnikov (2000, Theorem 1), it is known that as $T\to+\infty$, $\mathrm{Var}\bigl(\#\{j\geq 1\mid\mathfrak{a}_{j}^{\beta=2}>-T\}\bigr)\sim C\ln(T)$, where the value [^9] Soshnikov (2000) stated a different value of $C$. The arithmetic error was noticed and corrected in (Landon and Sosoe, 2022, Theorem 6.2). of the constant is $C=\frac{3}{4\pi^{2}}$. Through (8.15) this implies
$\lim_{T\to+\infty}\frac{\mathrm{Var}\bigl(\#\{j\geq 1\mid\mathfrak{a}_{j}>-T\}+\#\{j\geq 1\mid\mathfrak{a}^{\prime}_{j}>-T\}\bigr)}{\ln(T)}=4\cdot C,$
because the “even” operator divides the number of particles by two (up to error of at most $1$), and therefore divides the variance by four (up to an error negligible as $T\to\infty$). Since variances for independent random variables are added, we get (8.14) with right-hand side $2C$. ∎
@@
For each $\varepsilon>0$ there exists a random variable $\mathfrak{J}=\mathfrak{J}(\varepsilon)$, such that
(8.16) $\left|\mathfrak{a}_{j}+\left(\frac{3\pi j}{2}\right)^{2/3}\right|\leq j^{\varepsilon},\qquad\text{ almost surely for all }j>\mathfrak{J}.$
@@
$\bigr|\#\{j\geq 1\mid\mathfrak{a}_{j}>-n\}-\mathbb{E}\#\{j\geq 1\mid\mathfrak{a}_{j}>-n\}\bigl|>n^{1/2+\varepsilon}.$
@@
$\bigr|\#\{j\geq 1\mid\mathfrak{a}_{j}>-n\}-\mathbb{E}\#\{j\geq 1\mid\mathfrak{a}_{j}>-n\}\bigl|\leq n^{1/2+\varepsilon},\qquad\text{for all }n>\mathfrak{n}.$
@@
$\left|\#\{j\geq 1\mid\mathfrak{a}_{j}>-n\}-\frac{2}{3\pi}n^{3/2}\right|\leq n^{1/2+\varepsilon},\qquad\text{for all }n>\mathfrak{n}.$
@@
Note that $\frac{1}{w-\mathfrak{a}_{j}}-\frac{1}{\mathbf{i}-\mathfrak{a}_{j}}=\frac{\mathbf{i}-w}{(w-\mathfrak{a}_{j})(\mathbf{i}-\mathfrak{a}_{j})}.$ Hence, using (8.16), the jj–th term in the first sum of (8.19) decays as $j^{-4/3}$ and the sum is absolutely convergent, uniformly in $w$ bounded away from $\mathfrak{a}_{j}$. For the second sum, its imaginary part is
@@
Using (8.16), for large $j$ we have $\mathfrak{a}_{j}=\left(\frac{3\pi j}{2}\right)^{2/3}+O(j^{\varepsilon})$, and, therefore, for large $m$ and $n$:
@@
We check the Cauchy criterion for (8.20) and compute the difference of its values at x=yx=-y and x=zx=-z for large $y>z>0$. Using (8.21), we get
@@
where $n(y)$ is the index $j$ for the closest to y-y point $\mathfrak{a}_{j}$ and $m(z)$ is the index for the closest to z-z point $\mathfrak{a}_{j}$. Assuming $y$, $z$ large, so that $n(y)>\mathfrak{J}$ and $m(z)>\mathfrak{J}$, we use (8.16) and get
@@
We condition on $\{\mathfrak{a}_{j}\}$ and note that we deal with a sum of independent mean 0 random variables. Hence, by the Kolmogorov two-series theorem (see, e.g., Durrett (2019, Theorem 2.5.6)), the almost sure convergence would follow from the convergence of the sum of (conditional) variances, i.e., convergence of the series
@@
$\lim_{x\to-\infty}\left[\sum_{j:\,\mathfrak{a}_{j}>x}\left(\frac{1}{w+\left(\frac{3\pi j}{2}\right)^{2/3}}+\frac{\xi_{j}^{2}-1}{w+\left(\frac{3\pi j}{2}\right)^{2/3}}+\xi_{j}^{2}\left[\frac{\left(\frac{3\pi j}{2}\right)^{2/3}+\mathfrak{a}_{j}}{(w-\mathfrak{a}_{j})(w+\left(\frac{3\pi j}{2}\right)^{2/3})}\right]\right)-\frac{2}{\pi}\sqrt{-x}\right].$
@@
We show that as $w\to\infty$ with $\mathrm{Re}(w)\geq 0$, the second and third sums in (8.24) vanish.